OpenStudy (anonymous):

Solve the following system of equations by linear combination: 3d – e = 7 d + e = 5 There is no solution. There are an infinite number of solutions. The solution is (3, 2). The solution is (2, –1).

OpenStudy (anonymous):

3d - e = 7 d + e = 5 ---------add 4d = 12 -- divide by 4 d = 3 now sub 3 in for d in either of the equations d + e = 5 3 + e = 5 e = 5 - 3 e = 2 solution is (3,2)

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

no, thanks you

OpenStudy (anonymous):

YW :)